OpenStudy (anonymous):

How many subsets are there for a set with 15 elements?

OpenStudy (amoodarya):

every set has 2^n subset n=number of elements

OpenStudy (anonymous):

What does that mean?

OpenStudy (anonymous):

\[2^15\] in othe words . 2*2*2*2*2*2*2*2*2*2*2*2*2*2*2 =1024*32 =32768

OpenStudy (anonymous):

What??? I'm lost -_-